# GLEE: A Unified Framework and Benchmark for Language-based Economic Environments

### v1

@@ Abstract
Large Language Models (LLMs) show significant potential in economic and strategic interactions, where communication via natural language is often prevalent. This raises key questions: Do LLMs behave rationally? Can they mimic human behavior? Do they tend to reach an efficient and fair outcome? What is the role of natural language in the strategic interaction? How do characteristics of the economic environment influence these dynamics? These questions become crucial concerning the economic and societal implications of integrating LLM-based agents into real-world data-driven systems, such as online retail platforms and recommender systems. While the ML community has been exploring the potential of LLMs in such multi-agent setups, varying assumptions, design choices and evaluation criteria across studies make it difficult to draw robust and meaningful conclusions. To address this, we introduce a benchmark for standardizing research on two-player, sequential, language-based games. Inspired by the economic literature, we define three base families of games with consistent parameterization, degrees of freedom and economic measures to evaluate agents’ performance (self-gain), as well as the game outcome (efficiency and fairness). We develop an open-source framework for interaction simulation and analysis, and utilize it to collect a dataset of LLM vs. LLM interactions across numerous game configurations and an additional dataset of human vs. LLM interactions. Through extensive experimentation, we demonstrate how our framework and dataset can be used to: (i) compare the behavior of LLM-based agents to human players in various economic contexts; (ii) evaluate agents in both individual and collective performance measures; and (iii) quantify the effect of the economic characteristics of the environments on the behavior of agents. We believe that our framework can contribute to the growing intersection of LLMs, ML, and economics, and we encourage researchers to explore it further and build on its foundation.
@@ 1 Introduction
Recent research has increasingly focused on the capabilities of Large Language Models (LLMs) in decision-making tasks, revealing their potential to operate as autonomous agents in various economic environments, that typically require involving complex strategic thinking (Horton 2023; Wang et al. 2023; Zhu et al. 2024; Li et al. 2024). Applications of LLM-based agents in such environments include Chess-playing (Feng et al. 2024), task-oriented dialogue handling (Ulmer et al. 2024), financial advisement (Lakkaraju et al. 2023) and beyond. The promising capabilities of LLMs in strategic decision-making call for the study of these agents’ behavior through the lens of game theory, e.g. whether and when LLM-based agents naturally converge to Nash-equilibrium (Guo et al. 2024), and how well they learn to cooperate in repeated interactions (Akata et al. 2023). The rise of LLM agents also calls for a clear benchmark for assessing the extent to which such agent behaves rationally, which is addressed by a recent work of Raman et al. (2024).
An important property of many real-world economic environments is that communication between agents typically occurs through natural language. While economic modeling usually abstracts away these nuances for the sake of keeping the model simple and tractable, in practical scenarios the fact that natural language is involved may significantly affect the interaction outcome. For instance, in bargaining between two parties, the same offer can be interpreted very differently depending on how it is framed. Consider the following two offers for a business partnership: (a) ”We propose a 40-60 split as we believe your expertise will drive the majority of the success.” (b) ”We propose a 40-60 split because we’ve already invested significant resources and need a lower share to maintain balance.” Both propose the same numerical terms, but the framing in each case may lead to different reactions based on how the rationale is presented. A similar rationale applies to a broader range of economic scenarios. To illustrate this further, let us consider the following representative use cases:
@@
  • Negotiation. Alice aims to sell a product to a potential buyer, Bob. They negotiate by taking turns proposing prices, with Alice deciding whether to accept Bob’s offer and sell, and Bob deciding whether to accept Alice’s price and buy, continuing this process until they reach an agreement or the negotiation concludes. The negotiation often involves natural language, such as descriptions of the product, which can influence the outcome. [^1] Notice that in the economic literature, the terms ”bargaining” and ”negotiation” are often used interchangeably to describe both the division of a divisible good and the process of price negotiation. In this paper, we use the terminology ”bargaining” to describe the former and ”negotiation” for the latter, for ease of exposition. We refer to Appendix A for further discussion on the differences between the two games.
@@
  • Persuasion. Alice is a seller trying to sell a product to Bob at an exogenously set fixed price. Alice knows the true quality of the product, while Bob only has a rough expectation of the product’s quality. Alice sends a message to persuade Bob to buy the product, aiming to convince him of its value, regardless of its true quality. Bob, however, benefits only if the product is genuinely high quality. The interaction then repeats for multiple rounds, where product qualities are realized independently at different rounds. In such a scenario, Alice balances between her maximizing one-shot gain and building positive reputation, and linguistic communication can play a significant role.
@@
These three types of interactions are inspired by influential models in the economics literature. They are also broad and flexible enough to capture a wide variety of real-world applications. The bargaining model, inspired by the seminal work of Rubinstein (1982), forms the basis for understanding how parties negotiate the division of shared resources, applicable in scenarios such as business mergers, partnership dissolution, and legal settlements. The negotiation game reflects celebrated models of bilateral trade (Myerson & Satterthwaite 1983; McAfee 1992) where buyers and sellers negotiate over the prices of an indivisible good, which often arises in various applications, including real estate transactions, corporate acquisitions, and e-commerce platforms. Lastly, the persuasion model draws on classical models of information asymmetry and strategic communication (Akerlof 1978; Crawford & Sobel 1982; Farrell & Rabin 1996), which play a pivotal role in advertising, marketing, political campaigning, and recommendation systems.
A key feature of all these games is that they are sequential, meaning players take turns acting rather than acting simultaneously. This makes communication, and in particular language-based communication, a crucial element of the interaction. In sequential games, communication is typically direct, with players able to attach messages to their actions (i.e., in the bargaining example, the textual message is attached to the proposal). In contrast, in simultaneous games, communication tends to be indirect, as players act independently and typically learn to cooperate by observing and reacting to past actions rather than through direct message exchange. A prime example of a simultaneous, language-based economic interaction is the competition among web publishers in search engines. Publishers simultaneously create content (usually in the form of textual documents) for their websites, and then gradually learn and adapt based on outcomes, such as their relative search engine rankings and exposure. However, they usually lack the opportunity for real-time communication or coordination during this process. These simultaneous, language-based interactions are well-studied within the information retrieval community, and the induced publishers’ game is indeed modeled as a simultaneous game. [^2] See e.g. Raifer et al. (2017); Ben-Porat et al. (2019); Kurland & Tennenholtz (2022); Hron et al. (2022); Madmon et al. (2023; 2024); Nachimovsky et al. (2024). In these works, SEO competitions are modeled as simultaneous games. Our work is complementary, focusing on sequential settings where direct, language-based communication is prevalent.
@@
The AI and NLP communities have recognized the potential of integrating natural language into stylized economic models of bargaining, negotiation and persuasion. They also recognize the importance of studying the capabilities and limitations of LLM-based agents within such frameworks. A growing body of research is dedicated to evaluating and optimizing LLM-based agents in various bargaining and negotiation scenarios (Abdelnabi et al. 2023; Deng et al. 2024; Xia et al. 2024; Bianchi et al. 2024; Noh & Chang 2024). In the realm of persuasion games, recent studies have introduced language-based frameworks that explore the design of optimal information transmission policies (Raifer et al. 2022) and develop methods for generating data to predict human players’ behavior (Apel et al. 2022; Shapira et al. 2023; 2024).
@@
Although each of these economic setups has been studied in the context of LLM-based agents, the approaches have varied widely across different studies. Each paper often adopts distinct modeling and implementation assumptions, poses unique research questions, and applies diverse evaluation criteria, making it challenging to compare results and draw broader conclusions about the capabilities and limitations of LLM-based agents in economic environments. To enhance the real-world reliability of LLM-based agents, it is essential to establish a clear benchmark that provides a standardized framework for modeling these economic interactions. This benchmark would ensure comparability across studies and enable generalization of findings, facilitating a deeper understanding of how various factors influence interaction outcomes and leading to more robust and reliable conclusions about the performance of LLM-based agents in real-life economic situations.
1.1 Our contribution
@@
We introduce GLEE, a unified framework for Games in Language-based Economic Environments, [^3] Code and data are available in https://github.com/eilamshapira/GLEE. focusing on the case of two-player games. [^4] The two-player case is particularly important since the main shift occurs when moving from single-agent decision-making to two-agent interactions, in which strategic behavior and communication play crucial roles. Two-player scenarios are also widely observed in real life, and the phenomena emerging from these interactions tend to be more intuitive and easier to analyze, while still capturing the core behaviors that arise in more complex multi-agent environments. Our framework relies on a clear and comprehensive parameterization of the space of all bargaining, negotiation and persuasion games (as described above), and defines degree of freedom and evaluation metrics that are consistent across economic contexts. The generality of the space of games spanned by our parameterization follows from the richness of the degrees of freedom considered, which include the game horizon (number of rounds within each game), information structure (whether agents know each others’ preferences or not), and communication form (whether agents communicate via free language of structured messages).
@@
Through this framework, we have collected a dataset of LLM vs. LLM games, containing 7.15M decisions made by LLM-based agents in over 954K games, using four different LLMs. Our framework allows for running controlled experiments across various degrees of freedom, language models, and game scenarios, making the dataset valuable for comparing and evaluating LLM performance in different economic setups. It also facilitates deeper insights into how characteristics of the economic environment influence agent behavior and ultimately shape the outcomes of interactions. Additionally, we design and develop an interactive interface we run experiments in which human players play against LLM-based agents in those considered games, for a variety of game parameters, and collect a human vs. LLM interaction dataset.
To recap, the contribution of this paper is fourfold. First, we introduce a clear parametrization of a large, general, and representative set of sequential, two-player, language-based games. This parametrization is inspired by both the relevant economic literature, as well as advances in the LLM agents literature. Importantly, our parameterization can be naturally extended beyond the two-player case. Second, we release open-source code for GLEE, a unified framework that enables controlled experiments across the representative families of games considered. Notably, our framework is flexible and facilitates the collection of additional interactions with novel LLMs, game parameters and evaluation metrics, which is crucial given the fast advancements in the field. Third, we utilize GLEE to collect a dataset of LLM vs. LLM and human vs. LLM interactions in a wide range of economic environments, building on the foundations of our comprehensive parametrization. Finally, we demonstrate how the dataset can be analyzed to draw meaningful conclusions and economic insights regarding LLM-based players and human players in language-based economic environments.
@@ 2 Game Families and Parametrization
In this section, we formally define the three game families discussed in the introduction: bargaining, negotiation and persuasion. [^6] In Appendix A we further discuss the related economic literature, and review some well-known theoretical results concerning these games. For each family, we provide a formal game-theoretic definition of the game (including the players, strategies and utilities), define degrees of freedom, and define the game outcome and evaluation metrics. While previous research has effectively tackled the challenge of evaluating the rationality of individual agents (Raman et al. 2024), our focus extends to evaluating the outcome reached by the strategic behavior of the agents, using the fundamental economic notions of efficiency and fairness.  [^5] Metrics are normalized to $[0,1]$, such that higher values indicate better (more efficient or fair) outcomes.
In our modeling, we focus on three main degrees of freedom that are critical for understanding the dynamics of LLM-based agents in economic interactions: game horizon, information structure, and communication form. The game horizon refers to the number of time periods during which the game is played and whether the length of the horizon is known or unknown to the agents. [^7] In economic theory, ”infinite horizon” typically refers to a large, unspecified duration. Accordingly, we use the term ”infinite horizon” to describe cases where the horizon is both large and unknown to the agents. This factor influences the strategies agents adopt, particularly in terms of long-term planning and anticipation of future moves. The information structure determines whether agents are aware of each other’s preferences, impacting their ability to predict and respond to the actions of others. Lastly, the communication form specifies whether communication between agents occurs through free language or structured, concise messages, which affects the richness and clarity of the exchanges.
@@ 2.1 Bargaining games
The first family of games is inspired by the celebrated bargaining model of Rubinstein (1982). The model encompasses a class of bargaining games where two players, Alice and Bob, alternate offers over a time horizon $T$ (usually, $T=\infty$) to divide a fixed sum of money $M$ between them. [^8] It is clear that under full rationality assumption, the amount of money does not play a role in the analysis. However, for human (or LLM) players, it is evident that the amount of money indeed matters (See Figures 16, 22, 28 in Appendix E). Importantly, in these games delays are costly, a concept captured by discount factors $\delta_{A},\delta_{B}$ assigned to each player, reflecting the decreasing value of future payoffs as time progresses. Formally:
@@
In our experiments, we simulate a wide range of such bargaining games, differing in the following degrees of freedom: (i) whether or not the players know their opponent’s discount factor (complete vs. incomplete information); (ii) whether or not the players know when the game terminates (finite vs. infinite); (iii) whether or not players communication involve natural language (structured vs. linguistic); (iv) the values of $\delta_{A},\delta_{B}\in(0,1)$ and $M$; (v) the value of $T$ in the finite horizon case.
@@ 2.2 Negotiation games
As in the case of bargaining games, Alice and Bob alternate offers: at each odd stage, Alice posts a price and Bob decides whether to buy the product or move on to the next stage. At each even stage, Bob is the one to post a price and Alice decides whether to sell at this price or reject and move on to the next stage. The game is played for $T$ stages, which again can be either finite or ”infinite” (i.e., large and unknown to both players). Unlike the bargaining game, here we assume no discount factors on the utilities, hence whenever a price $p$ is accepted, the utilities for Alice and Bob are $p-V_{A}$ and $V_{B}-p$ respectively. If no trade is made, then the utilities are defined to be $(0,0)$.
In this class of games, we consider the following degrees of freedom: (i) whether or not the players know their opponent’s product valuation (complete vs. incomplete information); (ii) whether or not the players know when the game terminates (finite vs. infinite); (iii) whether or not players communication involve natural language (structured vs. linguistic); (iv) the values of $F_{A},F_{B}\in(0,1)$ and $M$; (v) the value of $T$ in the finite horizon case.
An outcome of a negotiation game is captured by $p_{ev}$, which is the price at which the product is sold when there is a trade, and defined to be $p_{ev}=\emptyset$ whenever no trade is made. We consider the following evaluation measures of the game outcome fairness and efficiency. When there is trade, fairness is measured by $1-4\cdot\left(\frac{p_{ev}-p_{f}}{M}\right)^{2}$, where $p_{f}=\frac{V_{A}+V_{B}}{2}$ is the ”fairest price”. [^10] Notice that unlike all other metrics, the fairness metric in negotiation games is not normalized, as the price $p_{ev}$ is unbounded in principle. However, we observe that normalizing the difference $p_{ev}-p_{f}$ by the scale parameter $M$ results in a measure that is between 00 and $1$ in 99.5% of the cases. When trade is not made, we define the fairness to be $1$ (i.e., maximal fairness) to reflect that no-trade does not change the default allocation of the product. Efficiency is defined to be $1$ in the following cases (and zero otherwise): (a) Alice values the product more than Bob, and does not sell it ($V_{A}\geq V_{B}$ and $p=\emptyset$); (b) Alice values the product less Bob, and sells the product at a price that is beneficial for both players ($V_{A}\leq p\leq V_{B}$); When averaged over a large number of simulated games for a certain game configuration, this measure estimated the probability of the event ”an efficient trade occurs when it should occur”.

### v2

@@ Abstract
Large Language Models (LLMs) show significant potential in economic and strategic interactions, where communication via natural language is often prevalent. This raises key questions: Do LLMs behave rationally? How do they perform compared to humans? Do they tend to reach an efficient and fair outcome? What is the role of natural language in strategic interaction? How do characteristics of the economic environment influence these dynamics? These questions become crucial concerning the economic and societal implications of integrating LLM-based agents into real-world data-driven systems, such as online retail platforms and recommender systems. While the ML community has been exploring the potential of LLMs in such multi-agent setups, varying assumptions, design choices and evaluation criteria across studies make it difficult to draw robust and meaningful conclusions. To address this, we introduce a benchmark for standardizing research on two-player, sequential, language-based games. Inspired by the economic literature, we define three base families of games with consistent parameterization, degrees of freedom and economic measures to evaluate agents’ performance (self-gain), as well as the game outcome (efficiency and fairness). We develop an open-source framework for interaction simulation and analysis, and utilize it to collect a dataset of LLM vs. LLM interactions across numerous game configurations and an additional dataset of human vs. LLM interactions. Through extensive experimentation, we demonstrate how our framework and dataset can be used to: (i) compare the behavior of LLM-based agents in various economic contexts; (ii) evaluate agents in both individual and collective performance measures; and (iii) quantify the effect of the economic characteristics of the environments on the behavior of agents. Our results suggest that the market parameters, as well as the choice of the LLMs, tend to have complex and interdependent effects on the economic outcome, which calls for careful design and analysis of the language-based economic ecosystem.
@@ 1 Introduction
Recent research has increasingly focused on the capabilities of Large Language Models (LLMs) in decision-making tasks, revealing their potential to operate as autonomous agents in various economic environments, that typically require involving complex strategic thinking [20, 47, 52, 28, 44]. Applications of LLM-based agents in such environments include Chess-playing [17], task-oriented dialogue handling [46], financial advisement [27] and beyond. The promising capabilities of LLMs in strategic decision-making call for the study of these agents’ behavior through the lens of game theory, e.g. whether and when LLM-based agents naturally converge to Nash-equilibrium [18], and how well they learn to cooperate in repeated interactions [2]. The rise of LLM agents also calls for a clear benchmark for assessing the extent to which such agents behave rationally [38].
An important property of many real-world economic environments is that communication between agents typically occurs through natural language. While economic modeling usually abstracts away these nuances for the sake of keeping the model simple and tractable, in practical scenarios the fact that natural language is involved may significantly affect the interaction outcome. For instance, in bargaining between two parties, the same offer can be interpreted very differently depending on how it is framed. Consider the following two offers for a business partnership: (a) "We propose a 40-60 split as we believe your expertise will drive the majority of the success." (b) "We propose a 40-60 split because we’ve already invested significant resources and need a lower share to maintain balance." Both offer the same numerical terms, but the framing in each case may lead to different reactions based on how the rationale is presented. A similar rationale applies to a broader range of economic scenarios. To illustrate this further, let us consider the following representative use cases:
@@
  • Negotiation. Alice aims to sell a product to a potential buyer, Bob. They negotiate by taking turns proposing prices, with Alice deciding whether to accept Bob’s offer and sell, and Bob deciding whether to accept Alice’s price and buy, continuing this process until they reach an agreement or the negotiation concludes. The negotiation often involves natural language, such as descriptions of the product, which can influence the outcome. [^1] Notice that in the economic literature, the terms ”bargaining” and ”negotiation” are often used interchangeably to describe both the division of a divisible good and the process of price negotiation. In this paper, we use the terminology ”bargaining” to describe the former and ”negotiation” for the latter, for ease of exposition. We refer to Appendix C for further discussion on the differences between the two games.
@@
  • Persuasion. Alice is a seller trying to sell a product to Bob at an exogenously set fixed price. Alice knows the true quality of the product, while Bob only has a rough expectation of the product’s quality. Alice sends a message to persuade Bob to buy the product, aiming to convince him of its value, regardless of its true quality. Bob, however, benefits only if the product is of genuinely high quality. The interaction then repeats for multiple rounds, with the product quality in each round being independently drawn. In such a scenario, Alice balances between her maximizing one-shot gain and building positive reputation, and linguistic communication can play a significant role.
@@
These three types of interactions are inspired by influential models in the economics literature. They are also broad and flexible enough to capture a wide variety of real-world applications. The bargaining model, inspired by the seminal work of Rubinstein [40], forms the basis for understanding how parties negotiate the division of shared resources, applicable in scenarios such as business mergers, partnership dissolution, and legal settlements. The negotiation game reflects celebrated models of bilateral trade [32, 31] where buyers and sellers negotiate over the prices of an indivisible good, which often arises in various applications, including real estate transactions, corporate acquisitions, and e-commerce platforms. Lastly, the persuasion model draws on classical models of information asymmetry and strategic communication [3, 12, 16], which play a pivotal role in advertising, marketing, political campaigning, and recommendation systems. [^2] These three fundamental game families are sequential games, in which players act in turns, unlike in simultaneous games. We discuss the importance of studying such games in the context of language-based environments in Appendix B.
@@
The AI and NLP communities have recognized the potential of integrating natural language into stylized economic models of bargaining, negotiation and persuasion. They also recognize the importance of studying the capabilities and limitations of LLM-based agents within such frameworks. Research increasingly focuses on evaluating and improving LLM agents in bargaining and negotiation settings [1, 13, 50, 9, 34, 22]. In the realm of persuasion games, recent studies have introduced language-based frameworks that explore the design of optimal information transmission policies [36] and develop methods for generating data to predict human players’ behavior [4, 42, 43].
@@
Although each of these economic setups has been studied in the context of LLM-based agents, the approaches have varied widely across different studies. Each paper often adopts distinct modeling and implementation assumptions, poses unique research questions, and applies diverse evaluation criteria, making it challenging to compare results and draw broader conclusions about the capabilities and limitations of LLM-based agents in economic environments. Moreover, the vast majority of the existing game-theoretic benchmarks for LLMs do not even include setups in which agents are allowed to communicate in natural language [14, 23, 18, 48].
To enhance the real-world reliability of LLM-based agents, it is essential to establish a clear benchmark that provides a standardized framework for modeling these economic interactions. This benchmark would ensure comparability across studies and enable generalization of findings, facilitating a deeper understanding of how various factors influence interaction outcomes and leading to more robust and reliable conclusions about the performance of LLM-based agents in real-life economic situations.
Our Contribution
@@
We introduce GLEE, a unified framework for Games in Language-based Economic Environments, [^3] Code and data are available in https://github.com/eilamshapira/GLEE. focusing on the case of two-player games. [^4] The two-player case is particularly important since the main shift occurs when moving from single-agent decision-making to two-agent interactions, in which strategic behavior and communication play crucial roles. Two-player scenarios are also widely observed in real life, and the phenomena emerging from these interactions tend to be more intuitive and easier to analyze, while still capturing the core behaviors that arise in more complex multi-agent environments. Our framework provides a principled way to evaluate the performance of LLM agents and human players in a wide class of fundamental language-based economic scenarios. Central to the framework is a clear and comprehensive parameterization of the space of all bargaining, negotiation and persuasion games (as described above). It defines degrees of freedom and evaluation metrics that are consistent across economic contexts. The generality of the space of games spanned by our parameterization follows from the richness of the degrees of freedom considered, which include the game horizon (number of rounds within each game), information structure (whether agents know each others’ preferences or not), and communication form (whether agents communicate via free language or structured messages). The framework is implemented as an open-source codebase that allows researchers to instantiate a wide range of economic games and evaluate LLM behavior within them. While the specific LLMs used in experiments may evolve or become outdated, the evaluation setup and metrics remain valid. A detailed comparison between GLEE and existing benchmarks is provided in Appendix A.
@@
Using this framework, we have collected a dataset of LLM vs. LLM games, comprising 587K decisions made by LLM-based agents across more than 80K games, involving 13 different LLMs. The framework enables controlled experimentation across a wide range of language models, game types, and experimental conditions, making the dataset a valuable resource for evaluating LLM behavior in diverse economic settings. It also supports in-depth analyses of how the parameters of the economic environment affect agent strategies and shape interaction outcomes. Additionally, we developed an interactive interface that enables human participants to compete against LLM-based agents across various game configurations. Using this setup, we conducted experiments and compiled a complementary dataset of human vs. LLM interactions.
Through extensive analysis of the collected data, we uncover meaningful behavioral and economic phenomena. First, we find that the economic outcome, measured by efficiency, fairness, and agent self-gain, is significantly shaped by market parameters, including the information structure, communication form, and interaction horizon. Second, we show that there are no absolute best-performing models in terms of any of the evaluation measures, and that the performance of one LLM strongly depends on its competitor’s choice of LLM. Finally, we compare LLMs to human participants and find that human behavior is markedly more extreme: in each setting, they either outperform all LLMs or fall behind them entirely, depending on the game and their assigned role. Together, these findings show how GLEE can be used not only as an infrastructure for research, but also as a tool for deriving concrete insights into economic reasoning and strategic behavior in language-based interactions.
@@ 2 Game Families and Parametrization
In this section, we formally define the three game families discussed in the introduction: bargaining, negotiation and persuasion. In Appendix C we further discuss the related economic literature, and review some well-known theoretical results concerning these games. For each family, we provide a formal game-theoretic definition of the game (including the players, strategies and utilities), define degrees of freedom, and define the game outcome and evaluation metrics. While previous research has effectively tackled the challenge of evaluating the rationality of individual agents [38], our focus extends to evaluating the outcome reached by the strategic behavior of the agents, using the fundamental economic notions of efficiency and fairness.  [^5] Metrics are normalized to $[0,1]$, such that higher values indicate better (more efficient or fair) outcomes.
In our modeling, we focus on three main market characteristics that are critical for understanding the dynamics of LLM-based agents in economic interactions: game horizon, information structure, and communication form. The game horizon refers to the number of time periods during which the game is played and whether the length of the horizon is known or unknown to the agents. [^6] In economic theory, ”infinite horizon” typically refers to a large, unspecified duration. Accordingly, we use the term ”infinite horizon” to describe cases where the horizon is both large and unknown to the agents. This factor influences the strategies agents adopt, particularly in terms of long-term planning and anticipation of future moves. The information structure determines whether agents are aware of each other’s preferences, impacting their ability to predict and respond to the actions of others. Lastly, the communication form specifies whether communication between agents occurs through free language or structured, concise messages, which affects the richness and clarity of the exchanges.
@@ 2.1 Bargaining Games
The first family of games is inspired by the celebrated bargaining model of Rubinstein [40]. The model encompasses a class of bargaining games where two players, Alice and Bob, alternate offers over a time horizon $T$ (usually, $T=\infty$) to divide a fixed sum of money $M$ between them. [^7] It is clear that under a full rationality assumption, the amount of money does not play a role in the analysis. However, for human (or LLM) players, it is evident that the amount of money indeed matters. Importantly, in these games delays are costly, a concept captured by discount factors $\delta_{A},\delta_{B}$ assigned to each player, reflecting the decreasing value of future payoffs as time progresses. Formally:
@@
In our experiments, we simulate a wide range of such bargaining games, differing in the following degrees of freedom: (i) whether or not the players know their opponent’s discount factor (complete vs. incomplete information); (ii) whether or not the players know when the game terminates (finite vs. infinite); (iii) whether or not players communication involve natural language (structured vs. linguistic); (iv) the values of $\delta_{A},\delta_{B}\in(0,1)$ and $M$; and (v) the value of $T$ in the finite horizon case.
@@ 2.2 Negotiation Games
As in the case of bargaining games, Alice and Bob alternate offers: at each odd stage, Alice posts a price and Bob decides whether to buy the product or move on to the next stage. At each even stage, Bob is the one to post a price and Alice decides whether to sell at this price or reject and move on to the next stage. The game is played for $T$ stages, which again can be either finite or "infinite" (i.e., large and unknown to both players). Unlike the bargaining game, here we assume no discount factors on the utilities, hence whenever a price $p$ is accepted, the utilities for Alice and Bob are $p-V_{A}$ and $V_{B}-p$ respectively. If no trade is made, then the utilities are defined to be $(0,0)$.
In this class of games, we consider the following degrees of freedom: (i) whether or not the players know their opponent’s product valuation (complete vs. incomplete information); (ii) whether or not the players know when the game terminates (finite vs. infinite); (iii) whether or not players communication involve natural language (structured vs. linguistic); (iv) the values of $F_{A},F_{B}\in(0,1)$ and $M$; and (v) the value of $T$ in the finite horizon case.
An outcome of a negotiation game is captured by $p_{ev}$, which is the price at which the product is sold when there is a trade, and defined to be $p_{ev}=\emptyset$ whenever no trade is made. We consider the following evaluation measures of the game outcome fairness and efficiency. When there is trade, fairness is measured by $1-4\cdot\left(\frac{p_{ev}-p_{f}}{M}\right)^{2}$, where $p_{f}=\frac{V_{A}+V_{B}}{2}$ is the "fairest price". [^9] Notice that unlike all other metrics, the fairness metric in negotiation games is not normalized, as the price $p_{ev}$ is unbounded in principle. However, we observe that normalizing the difference $p_{ev}-p_{f}$ by the scale parameter $M$ results in a measure that is between 00 and $1$ in 99.5% of the cases. When trade is not made, we define the fairness to be $1$ (i.e., maximal fairness) to reflect that no-trade does not change the default allocation of the product. Efficiency is defined to be $1$ in the following cases (and zero otherwise): (a) Alice values the product more than Bob, and does not sell it ($V_{A}\geq V_{B}$ and $p=\emptyset$); and (b) Alice values the product less Bob, and sells the product at a price that is beneficial for both players ($V_{A}\leq p\leq V_{B}$); When averaged over a large number of simulated games for a certain game configuration, this measure estimated the probability of the event "an efficient trade occurs when it should occur".